OpenStudy (anonymous):

find all the roots of the equation x^3-3x^2+x-3=0

OpenStudy (anonymous):

(x^3 - 3x^2) + (x-3) = 0 x^2(x - 3) + (x - 3) = 0 (x^2 + 1) (x - 3) = 0

OpenStudy (anonymous):

\[x^{3}-3x ^{2}+x-3=0\]

OpenStudy (anonymous):

i need all the roots of that

OpenStudy (lollylau):

use moneybird's factored expression. either x^2+1 is zero or x-3 is 0.

OpenStudy (lollylau):

-1, 1, 3

OpenStudy (anonymous):

x^2 + 1 = 0 x^2 = -1 no solution from this x - 3 = 0 x = 3

OpenStudy (lollylau):

wait only one real solution?

OpenStudy (lollylau):

3

OpenStudy (anonymous):

yeah you can't square root negative number. You get imaginary solutions as well

OpenStudy (anonymous):

i dont understnd ineed step by step

OpenStudy (lollylau):

you factor the expression. you find the x's that make the expression zero